An unbalanced force of 2.5 x 10^3 N is applied to an object with a weight of 3.0 x 10^3 N on a frictionless surface. What is the object's acceleration?
1
Expert's answer
2019-03-05T09:42:35-0500

We can find the acceleration of the object from the Newton's Second Law of Motion:


"F = ma,"

here, "F" is the unbalanced force applied to the object, "m = \dfrac{W}{g}" is the mass of the object, "W" is the weight of the object, "g" is the acceleration due to gravity and "a" is the acceleration of the object.

Then, we get:


"a = \dfrac{F}{m} = \dfrac{Fg}{W} = \dfrac{2.5 \cdot 10^3 N \cdot 9.8 \dfrac{m}{s^2}}{3.0 \cdot 10^3 N} = 8.16 \dfrac{m}{s^2}."

Answer:

"a = 8.16 \dfrac{m}{s^2}."